Investment analysis q 2,30

Q12.30 As a manufacturer, you have to decide how ributors to sign up. Serv- many regional dist ing a distributor costs more the farther away it is from the factory, and different distributors have different demand. By region, gross rev enues and costs are in millions of dollars) as follows: Distributor A B CDEFG Gross Revenuc $5 SA SA S3 S2 $7 $1 Cost There is no lime value of money in this problem. (a) Is it feasible to work out all possible combinations of distributors service? Is it sensible? you can (b) Which regions should you deliver to? (c) What is the total profit for serving them? (d) What is the marginal benefit and cost of serving the least prolitable of your serviced distributors? (e) What would be the marginal benefit and cost of serving one more distributor? (f) Now assume that to get into this business, you would also have to set up the factory This would cost you a one-time upfront expense of $5 million. You can think of this as spreading the cost across distributors. How would this change your decision?
Answer #1

Distributor A- Gross profit = $5million-$2million= $3million

Distributor B- GP= $4million-$2million=$2million

Distributor C- GP= $3million-$4million=-$1million(Net Loss)

Distributor D- GP= $2million-$4million=-$2million(Net Loss)

Distributor E- GP= $7million-$5million=$2million

Distributor F- GP= $1million-$6million=-$5million(Net Loss)

1. The distributors serving which serves as a net loss for the manufacturer should not be served. All the distribitors which give positive gross profit should be served.

2. The regions served should be - A,B, E

3. Total Gross Profit= $3million + $2 Million +$2 Million = $7 Million

4. The least Profitable is Distributor E as the Gross Profit/Cost ratio is the lowest in its case. The marginal benefit = $2 million and marginal cost =$5million

5. Based on same concept as above

6. As this fee is a cost that the manufacturer cannot avoid if he wants to get into business, he should include it into his gross profit calculation.

Gross Profit considering only 1 year operation= $ 7million - $ 5 million= $ 2 million----->lower than before

But this cost would be capitalized across years of operation so this cannot be directly taken into calculation